OpenStudy (anonymous):

Find the total area and volume for the cylinder with the given measurement. r = 1, h = 1 The number of sq units of the total area exceeds the number of cu. units in the volume by a) 3(pi) b) 2(pi) c) (pi)

OpenStudy (theviper):

Total area of a cylinder:- \[2\pi r(r+h)\]

OpenStudy (theviper):

Therefore, total area is 2*pi*1(1+1) 2*pi*2 4pi

OpenStudy (theviper):

& volume of a cylinder:- \[\pi r^2h\]

OpenStudy (theviper):

=\[\pi \times1^2\times1=\pi 1\]

OpenStudy (theviper):

=pi

OpenStudy (anonymous):

Thank You!(:

OpenStudy (theviper):

U r welcome d^_^b